OpenStudy (anonymous):

which number should come next ? 145,122,101,82,65,

OpenStudy (anonymous):

Pattern recognition is an intuitive process. 145 - 122 = 23. 122 - 101 = 21 101 - 82 = 19 82 - 65 = 17 The sequence is \(a_0 = 145\); \(a_n = 122 - 2n\). So basically, we subtract 15 from 65 to get the correct answer.

OpenStudy (charlotte123):

^

OpenStudy (charlotte123):

Thus - The correct number that should come next would be 50.

OpenStudy (charlotte123):

@deraysha23 And welcome to OpenStudy! :)

OpenStudy (anonymous):

50

OpenStudy (nurali):

i think 50 145-122=23 122-101=21 101-82=19 82-65=17 65-50=15

OpenStudy (nurali):

Welcome to Openstudy @deraysha23